# Quasi-Deterministic Modeling of Sub-THz Band Access Channels in Street Canyon Environments

### v1

@@ Abstract
Sub-terahertz (sub-THz) frequencies (100–300 GHz) are expected to play a key role in beyond-5G and 6G mobile networks. However, their quasi-optical propagation characteristics require new channel models beyond sub-100 GHz extrapolations. This paper presents an extensive double-directional (D-D) channel measurement campaign conducted in an outdoor street-canyon environment at 154 GHz and 300 GHz under both line-of-sight (LoS) and non-line-of-sight (NLoS) conditions using an in-house-developed channel sounder. Based on these measurements, clustering with merged datasets across the two frequencies enables comparative analyses that identify both common and distinct multipath clusters, as well as the frequency dependence of cluster-level characteristics. A quasi-deterministic (QD) channel model is then proposed, combining deterministic components, such as LoS and single-bounce reflections from side walls, with random components. Large-scale parameters (path loss, delay spread, angular spread, and Rician K-factor) are also evaluated. These results provide valuable insights into sub-THz propagation in urban street canyons and contribute toward the development of accurate, channel models for future 6G systems.
Index Terms:
Sub-terahertz, dual-band channel sounding, outdoor multipath propagation, Quasi-deterministic channel model, street canyon, outdoor hot spot, multipath interactionI Introduction
Sixth-generation (6G) mobile communications are expected to achieve ultra-high data rates exceeding Gbps and ultra-low latency on the order of s, far surpassing the capabilities of current networks. To meet these stringent requirements, research has increasingly focused on the low-terahertz (low-THz) or sub-THz spectrum (– GHz), which offers contiguous channel bandwidths of several tens of GHz. A significant milestone was reached at the World Radiocommunication Conference 2023 (WRC-23), where it was agreed to consider allocating the – GHz, – GHz, – GHz, – GHz, and – GHz bands for future International Mobile Telecommunications (IMT) as part of the preliminary agenda for WRC-31 [1]. This regulatory progress, together with growing research in channel measurements, propagation modeling, and early standardization discussions in organizations such as 3GPP [2] and ETSI [3], has further strengthened expectations that sub-THz frequencies will become a cornerstone of mobile networks beyond 2030.
@@
Sub-THz hot spots in street environments provide ultra-high-capacity access in dense urban areas, as illustrated in Fig. 1, enabling transformative applications such as immersive augmented reality (AR), virtual reality (VR), mixed reality (MR), connected mobility, and real-time edge computing. Several measurement campaigns have recently investigated sub-THz propagation in urban street canyon environments to support future 6G channel modeling. Angle-resolved measurements at and GHz in a UMi street canyon were conducted, highlighting the sparse multipath structure of THz channels [6]. Similar results were obtained in double-directional (D-D) measurement campaigns at 145 GHz, which revealed dominant LoS and facade reflections with limited angular spread [7]. At 140 and 220 GHz, LoS path loss was found to approximate free-space predictions, while NLoS conditions suffered significantly higher attenuation due to building obstructions. Multipath richness was still observed, with notable contributions from facades, vehicles, and street-level scatterers [8]. By contrast, 306–371 GHz measurements in a campus street environment revealed much sparser propagation: wall reflections incurred 5–30 dB additional loss, and both delay and angular spreads were substantially smaller than those reported in sub-100 GHz 3GPP street-canyon models. Collectively, these results highlight a clear trend: as frequency increases from 140 to 370 GHz, multipath richness diminishes and channels become increasingly sparse, underscoring the need for new channel models tailored to THz propagation [9]. Additional large-scale efforts by collaborative studies in ETSI and ITU-R have provided datasets [10, 11] that are now feeding into early sub-THz channel model development for 3GPP. Collectively, these studies demonstrate that street hot spots are crucial deployment scenarios for THz access, where LoS dominance and a few strong reflections can be effectively exploited for reliable high-capacity connectivity.
@@
In this context, the present study reports an extensive D-D channel measurement campaign conducted in an outdoor street-canyon environment at 154 and GHz under both LoS and NLoS conditions using an in-house-developed channel sounder. Based on these measurements, clustering with merged datasets from the two frequency bands is performed, enabling comparative analyses that identify both common and distinct multipath clusters, as well as the frequency dependence of cluster-level characteristics. A quasi-deterministic (QD) channel model is then proposed, in which the deterministic components include the LoS path and single-bounce reflections from the two side walls, while the randomness of multipath components is modeled using an approach adapted from corridor scenarios [13] owing to their structural similarity. The deterministic components also incorporate randomness introduced by reflector’s surface conditions, represented by a two-state Markov process for presence and absence. Finally, the propagation channel is comprehensively characterized in terms of large-scale parameters (LSPs), including path loss (PL), delay spread (DS), azimuth angular spread (AS), and Rician K-factor (KF). The main contributions of this work are summarized as:
@@
Development of a QD channel model with deterministic components and random components modeled using a corridor-inspired approach; and
Comprehensive evaluation of large-scale parameters, namely PL, DS, AS, and Rician KF.
The remainder of this paper is organized into five sections. Section II outlines the measurement campaign, including scenario descriptions and a brief overview of the channel sounding setup. Section III details the post-processing procedures, such as dual-band joint noise filtering and clustering. Section IV examines the dominant propagation mechanisms by analyzing the trajectories of measured power spectra and validating them through ray-tracing. Section V presents the proposed QD channel model, while Section VI characterizes the channel in terms of path loss (PL) and large-scale parameters (LSPs). Finally, Section VII summarizes the key findings and highlights directions for future research.
@@ II-A Measurement Scenarios
The D-D channel measurements were conducted in a street canyon with a width of 20–25 m on the Niigata University campus. High-rise buildings (height m) line both sides of the street canyon, as illustrated in Fig. 2(a). In addition, there are large glass windows on the side walls. During the measurement, the Rx was placed in 30 distinct positions (Rx1–Rx30) at a height of m from the ground, emulating a mobile terminal. The Tx was fixed at a height of m above the ground and sequentially positioned at two locations at one end of the street canyon. One location provided a LoS scenario, with Tx1 corresponding to receiver positions Rx1–Rx23 (distance ranging from 10–80 m). The other location was positioned behind the corner of a building, blocking the LoS path to the receiver and creating an NLoS scenario, with Tx2 corresponding to receiver positions Rx24–Rx30 (distance ranging from 10–40 m).
Directional scanning measurement was conducted, and the setup is shown in Fig. 2(b). On both sides, dBi gain directional horn antennas with the narrow HPBWs of and were employed for both bands. The full azimuth (Az) sweep was conducted at both Rx and Tx, whereas the Rx and Tx antenna elevation (El) angles were fixed at and , respectively ( tilted). As a result, angular samples in the azimuth at both Tx and Rx were obtained (1,600 combinations). The antenna gains and radiation pattern effects were removed from the measured data in the post-processing.
@@ II-B Channel Sounding
Measurements were conducted using a GHz dual-band channel sounder developed from our earlier work [14]. Since the 300-GHz system has already been described in detail in [14], this paper highlights the dual-band extensions and their distinctive features. The system parameters are summarized in Table I.
@@
Polarization | Vertical |
To compensate for the effects of system response, a back-to-back (B2B) calibration was performed using waveguide attenuators (ATT) prior to the campaign [14]. The effect of the antenna gain was removed through over-the-air (OTA) calibration by aligning the antennas, and the tilting effect was also compensated for by post-processing. Furthermore, considering the propagation loss associated with the distance between the Tx and Rx, the signal-to-noise ratio (SNR) of the received signal in each band is expected to degrade significantly. This limitation restricts the achievable measurement range. To mitigate this issue, the continuously received sounding symbols are coherently averaged over multiple symbol durations, thereby improving the SNR and extending the effective measurement range.
@@ III-B Noise filtering for dual-band datasets
and is a margin ensuring no false detection is expected when testing the samples times the total number . Here, was chosen.
The common threshold in (7), applied to all samples in a single dataset, generally yields satisfactory performance. However, sample sets at two frequencies, even when processed under identical noise-filtering conditions, often produce CIRs that do not align well due to frequency-dependent behavior or false detections. This suggests that noise filtering can be enhanced by cross-checking the two datasets, leveraging the fact that propagation mechanisms remain fundamentally consistent across frequencies. In this approach, the threshold for each frequency is adaptively adjusted along the delay axis by increasing or decreasing the value from the common threshold at delay bins where samples should be retained or removed, respectively. The detailed procedure is provided in Algorithm 1, where , denoting the PDP value of at the delay bin corresponding to , is compared with the common threshold of slightly reduced by , i.e., . If this value exceeds the reduced threshold, the individual threshold at is updated as to revive a miss-removed path; otherwise, the individual threshold at is updated as to eliminate a miss-detected path.
Fig. 3 shows an example of clustering (Rx19), where all CIR samples are plotted along the delay axis, with the decision margin empirically set to dB. The results demonstrate that multipath components are consistently detected across the two frequencies. It is noted that the resulting dynamic range (DR), defined as the difference between the peak power and the noise floor (), ranges from approximately to dB, while the spurious-free dynamic range (SFDR), defined as the difference between the peak power and the threshold level, ranges from to dB in the LoS scenario.
@@ IV-A Dominant Scattering Process
Figs. 6 depicts the spatial variation of the PDPs along the Rx trajectory for both scenarios. Specifically, Fig. 6(a) shows the relative PDPs, where the delay is normalized to the LoS delay, while Fig. 6(b) presents the normal PDPs as a function of the absolute delay. At both frequencies, the propagation characteristics exhibit pronounced sparsity and simplicity. In the LoS scenario, as shown in Fig. 6(a), two major scattering processes that vary consecutively along the Rx trajectory are identified as SB reflections from the north wall (NW) and south wall (SW), with their trajectories validated using the image method ray-tracing applied to the simplified model in Fig. 5(b). Aside from the three main components, such as the LoS path and the two SB reflections, only a few additional comonents are observed, mainly originating from double-bounce (DB) reflections and other unidentified interactions with random objects. This phenomenon become more pronounced at GHz. In contrast, in the NLoS scenario, as shown in Fig. 6(b), two dominant propagation mechanisms that evolve consecutively along the Rx trajectory are approximately identified as the SB reflection from the SW and the DB reflection resulting from the SW-NW interaction.
The behavior of the extracted MPCs is illustrated in Figs. 7(a), 7(b), and 7(c), which present the variations in MPC distribution across the delay, AoD, and AoA domains, respectively, at both frequencies. In the LoS scenario, the SB reflections from the NW and SW dominate, whereas DB reflections are rarely observed, a trend that was further confirmed by the clustering results.
@@ IV-B Cluster Characteristics
Fig. 8(b) shows the path gain (PG) distribution of the clusters for each frequency band. The solid lines represent the LoS cluster PG across all Rx points in the LoS scenario, while the dashed lines denote the total PG of the NLoS clusters. For the LoS clusters, the average PG values are dB at GHz and dB at GHz, corresponding to a power difference of dB. This difference is dB smaller than the additional FSPL of dB, indicating that the GHz band experiences slightly greater degradation. However, for the NLoS clusters, the average PG values are dB at GHz and dB at GHz, yielding a power difference of dB, which is dB larger than the additional FSPL.
@@
V Channel modeling
Based on the investigation of the propagation mechanisms, this section proposes a QD channel model inspired by those developed for corridor scenarios [13]. The discussion begins with the mathematical representation of the channel, followed by the modeling and generation of the deterministic and random components that constitute the channel. It should be noted that the channel model is developed only for the LoS scenario, due to the limited amount of available data.
V-A QD channel model
Due to the short wavelength of sub-THz waves, specular reflection dominates and can be effectively predicted using deterministic methods such as ray tracing. Building on this characteristic as the deterministic component, a QD channel model[5, 12] further accounts for non-negligible interactions arising from higher-order reflections and scattering caused by object shape and surface roughness, modeling them as random components. Namely, the propagation channel can be modeled as
where the deterministic components, , and random components, are separately represented.
where ‘’ and ‘’ denote the single-bounce reflections by NW and SW, respectively. denotes a random variable that captures the randomness of the path amplitude, including its presence or absence.
Freq. | Transision probability matrix |
GHz | , |
GHz | , |
V-B Deterministic components
Fig. 9(a) illustrates the distribution of the differences between the measured and calculated powers for the SB reflections from the NW and SW at both frequencies, where missing Rx positions indicate the absence of a signal. Fig.9(b) further presents the empirical CDFs together with their fitted models. To capture this probabilistic behavior induced by factors such as surface roughness, these models can be combined with a two-state Markov chain to represent signal presence or absence, as proposed in [13]. This property of the Markov chain is represented by the transition probability matrix . The modeling of path appearance is carried out separately for the SB reflections from the NW and SW at each frequency. Table II summarizes the estimated transition probability matrices for each reflector.
Freq. | 1/ | |||
GHz | ||||
GHz |
V-C Random components
As described above, non-negligible random interactions should be incorporated into the channel model for a more accurate representation. Fig. 10 shows the distributions of relative PG for the deterministic and random components, where the SB reflections from the NW and SW are treated as deterministic components and all other components are treated as random. Evidently, the random components make a significant contribution to the overall power. As explicated earlier, MPCs originating from non-predominant higher-order reflections and difficult-to-identify propagation paths are modeled as random components.
and has been fitted accordingly in Fig. 11(a). Here, represents the arrival time of the th MPCs, and is the reciprocal of the mean interarrival time (arrival rate). The fitting results presented in Table III show that the arrival rate at GHz is ns, which is higher than ns at GHz. It should be noted that these values are more than an order of magnitude larger than those observed in indoor corridor environments, owing to the outdoor setting.
where , represents the slope, and represents the intercept. Furthermore, and represent the delay and path gain of the LoS component at distance , respectively. denotes a zero-mean Gaussian random variable with variance , representing shadow fading. The parameters and are estimated by fitting the plot in Fig.11(b) using the maximum likelihood method [19], considering a noise level of dB. The results yield and at GHz, and and at GHz, illustrating a similar trend at both frequencies. Furthermore, as shown in Fig. 11(c), the path gain deviation to the model expressed in dB fits a normal distribution with a standard deviations of at GHz and at GHz, respectively. The obtained parameters for random cluster model are tabulated in Table III.
Scen. | Freq. | CI Model | FI Model |
LoS | GHz | ||
GHz | |||
NLoS | GHz | ||
GHz |
VI Path Loss and LSP Characterization
Regardless of the specific channel model, channel parameters including PL and LSPs are essential for applying standard channel models such as 3GPP TR 38.901 [2]. In this section, the measured channels are characterized in terms of PL, DS, AS, and Rician KF at both frequencies and in both scenarios.
VI-A Path Loss

### v2

@@ Abstract
Sub-terahertz (sub-THz) frequencies (100–300 GHz) are expected to play a key role in beyond-5G and 6G mobile networks. However, their quasi-optical propagation characteristics require new channel models beyond sub-100 GHz extrapolations. This paper presents an extensive double-directional (D-D) channel measurement campaign conducted in an outdoor street-canyon environment at 154 GHz and 300 GHz under both line-of-sight (LoS) and non-line-of-sight (NLoS) conditions using an in-house-developed multi-tone frequency-domain channel sounder. Based on these measurements, clustering with merged datasets across the two frequencies enables comparative analyses that identify both common and distinct multipath clusters, as well as the frequency dependence of cluster-level characteristics. A quasi-deterministic (Q-D) channel model is then proposed, combining deterministic components, such as LoS and single-bounce reflections from side walls, with random components. Large-scale parameters (path loss, delay spread, angular spread, and Rician -factor) are also evaluated. These results provide valuable insights into sub-THz propagation in urban street canyons and contribute toward the development of accurate, channel models for future 6G systems.
I Introduction
Sixth-generation (6G) mobile communications are expected to achieve ultra-high data rates exceeding Gbps and ultra-low latency on the order of s, far surpassing the capabilities of current networks. To meet these stringent requirements, research has increasingly focused on the low-terahertz (low-THz) or sub-THz spectrum (– GHz)111Throughout this paper, “sub-THz” (or “low-THz”) refers to the – GHz range; unless otherwise stated, all results and model parameters are derived from measurements at GHz and GHz only (i.e., no extrapolation beyond GHz is performed), while “THz” is used only in a general sense to denote frequencies THz. , which offers contiguous channel bandwidths of several tens of GHz. A significant milestone was reached at the World Radiocommunication Conference 2023 (WRC-23), where it was agreed to consider allocating the – GHz, – GHz, – GHz, – GHz, and – GHz bands for future International Mobile Telecommunications (IMT) as part of the preliminary agenda for WRC-31 [1]. This regulatory progress, together with growing research in channel measurements, propagation modeling, and early standardization discussions in organizations such as 3GPP [2] and ETSI [3], has further strengthened expectations that sub-THz frequencies will become a cornerstone of mobile networks beyond 2030.
@@ I Introduction
THz hot spots in street environments provide ultra-high-capacity access in dense urban areas, as illustrated in Fig. 1, enabling transformative applications such as immersive augmented reality (AR), virtual reality (VR), mixed reality (MR), connected mobility, and real-time edge computing. Several measurement campaigns have recently investigated sub-THz propagation in urban street canyon environments to support future 6G channel modeling. Angle-resolved measurements at and GHz in a UMi street canyon were conducted, highlighting the sparse multipath structure of sub-THz channels [6]. Similar results were obtained in double-directional (D-D) measurement campaigns at 145 GHz, which revealed dominant LoS and facade reflections with limited angular spread [7]. At 140 and 220 GHz, LoS path loss was found to approximate free-space predictions, while NLoS conditions suffered significantly higher attenuation due to building obstructions. Multipath richness was still observed, with notable contributions from facades, vehicles, and street-level scatterers [8]. By contrast, 306–371 GHz measurements in a campus street environment revealed much sparser propagation: wall reflections incurred 5–30 dB additional loss, and both delay and angular spreads were substantially smaller than those reported in sub-100 GHz 3GPP street-canyon models. Collectively, these results highlight a clear trend: as frequency increases from 140 to 370 GHz, multipath richness diminishes and channels become increasingly sparse, underscoring the need for new channel models tailored to propagation in this frequency range [9]. Additional large-scale efforts by collaborative studies in ETSI and ITU-R have provided datasets [10, 11] that are now feeding into early sub-THz channel model development for 3GPP. Collectively, these studies demonstrate that street hot spots are crucial deployment scenarios for THz access, where LoS dominance and a few strong reflections can be effectively exploited for reliable high-capacity connectivity.
@@
In this context, the present study reports an extensive D-D channel measurement campaign conducted in an outdoor street-canyon environment at 154 and GHz under both LoS and NLoS conditions using an in-house-developed channel sounder. Based on these measurements, clustering with merged datasets from the two frequency bands is performed, enabling comparative analyses that identify both common and distinct multipath clusters, as well as the frequency dependence of cluster-level characteristics. A quasi-deterministic (Q-D) channel model is then proposed, in which the deterministic components include the LoS path and single-bounce reflections from the two side walls, while the randomness of multipath components is modeled using an approach adapted from corridor scenarios [13] owing to their structural similarity (e.g., “canyon-like” environments). The deterministic components also incorporate randomness introduced by reflector’s surface conditions, represented by a two-state Markov process for presence and absence. Finally, the propagation channel is comprehensively characterized in terms of path loss (PL), modeled using both Close-In (CI) and Floating-Intercept (FI) approaches. Additionally, we analyze large-scale parameters (LSPs), including root-mean-square (RMS) delay spread (DS), azimuth angular spread (AS), and Rician -factor (KF). The main contributions of this work are summarized as:
@@
Comprehensive evaluation of large-scale parameters, namely PL, DS, AS, and Rician -factor;
Development of a Q-D channel model that extends the corridor-based framework [13] to outdoor street canyons by integrating deterministic and random components.
The remainder of this paper is organized as follows. Section II outlines the measurement campaign, including scenario descriptions and a brief overview of the channel sounding setup. Section III details the post-processing procedures, such as dual-band joint noise filtering and clustering. Section IV examines the dominant propagation mechanisms by analyzing the trajectories of measured power spectra and validating them through ray-tracing. Section V characterizes the channel in terms of path loss (PL) and large-scale parameters (LSPs). Section VI presents the proposed Q-D channel model and its quantitative validation. Finally, Section VII summarizes the key findings and highlights directions for future research.
@@ II-A Measurement Scenarios
The D-D channel measurements were conducted in a street canyon with a width of 20–25 m on the Niigata University campus. High-rise buildings (height m) line both sides of the street canyon, as illustrated in Fig. 2(a). In addition, there are large glass windows on the side walls. During the measurement, the Rx was placed in 30 distinct positions (Rx1–Rx30) at a height of m from the ground, emulating a user equipment (UE). The Tx was fixed at a height of m above the ground to emulate a base station (BS) and sequentially positioned at two locations at one end of the street canyon. One location provided a LoS scenario, with Tx1 corresponding to receiver positions Rx1–Rx23 (distance ranging from 10–80 m). The other location was positioned behind the corner of a building, blocking the LoS path to the receiver and creating an NLoS scenario, with Tx2 corresponding to receiver positions Rx24–Rx30 (distance ranging from 10–40 m).
Directional scanning measurement was conducted, and the setup is shown in Fig. 2(b). On both sides, dBi gain directional horn antennas with the narrow HPBWs of and were employed for both bands. The full azimuth (Az) sweep was conducted at both Rx and Tx, whereas the Rx and Tx antenna elevation (El) angles were fixed at and , respectively ( tilted). Owing to the simple geometry of the environment, no elevation sweep was conducted [14, 15]; thus, the measurements focused on propagation characteristics in the local horizontal plane. As a result, angular samples in the azimuth at both Tx and Rx were obtained (1,600 combinations). The antenna gains and radiation pattern effects were removed from the measured data in the post-processing.
@@ II-B Channel Sounding
Measurements were conducted using a GHz dual-band multi-tone frequency-domain channel sounder developed from our earlier work [15]. Since the 300-GHz system has already been described in detail in [15], this paper highlights the dual-band extensions and their distinctive features. The system parameters are summarized in Table I.
@@
| Polarization | Vertical | |||||||
To compensate for the effects of system response, a back-to-back (B2B) calibration was performed using waveguide attenuators (ATT) prior to the campaign [15]. The antenna gain was de-embedded through over-the-air (OTA) calibration, where the antennas were boresight-aligned to maximize the received power, and the fixed elevation tilt was compensated for during post-processing using the antenna radiation patterns222The elevation tilt modifies effective antenna gain rather than intrinsic propagation mechanisms. With narrow-beam antennas, this significantly alters coupling to dominant paths. Given the fixed tilt, we focused on the local horizontal plane, compensating for antenna pattern effects in post-processing.. Furthermore, considering the propagation loss associated with the distance between the Tx and Rx, the signal-to-noise ratio (SNR) of the received signal in each band is expected to degrade significantly. This limitation restricts the achievable measurement range. To mitigate this issue, the continuously received sounding symbols are coherently averaged over multiple symbol durations, thereby improving the SNR and extending the effective measurement range.
@@ III-B Noise filtering for dual-band datasets
and is a margin designed to effectively suppress false detections is expected when testing the samples times the total number . Here, was chosen.
The common threshold in (7), applied to all samples in a single dataset, generally yields satisfactory performance. However, sample sets at two frequencies, even when processed under identical noise-filtering conditions, often produce CIRs that do not align well due to frequency-dependent behavior or false detections. This suggests that noise filtering can be enhanced by cross-checking the two datasets, leveraging the fact that propagation mechanisms remain fundamentally consistent across frequencies. In this approach, the threshold for each frequency is adaptively adjusted along the delay axis by exploiting the physical consistency of multipath propagation across the two bands. This process is designed to mitigate the DR imbalance caused by frequency-dependent losses (e.g., surface roughness). The detailed procedure, summarized in Algorithm 1, involves cross-referencing both power levels and angle-domain consistency.
For a unique candidate path detected in band at delay , the corresponding response in the complementary band is examined to distinguish between two scenarios:
1.
Miss-Detection (Rescue): First, the algorithm checks if a valid signal exists in that was suppressed by the common threshold. If the maximum power in at delay , denoted as , falls within a specific rescue margin below the common threshold (i.e., ) and the estimated angles in both bands are spatially aligned, the absence of the path in is attributed to a miss-detection. Consequently, the individual threshold for is locally decreased as to rescue the path.
2.
False Alarm (Removal): Conversely, if the component in fails this validation, either because its power is below the rescue margin or because the spatial angles do not match the candidate in , the peak in is regarded as a noise artifact (false alarm). In this case, the individual threshold for is locally increased as to remove this noise component.
This bidirectional validation ensures that weak high-frequency components are preserved only when spatially supported by the lower frequency, while simultaneously preventing the higher-DR band from introducing spurious noise peaks. This strategy effectively mitigates bias favoring the band with a larger dynamic range (154 GHz), particularly when valid components at 300 GHz are significantly attenuated by interaction loss. By enforcing strict angle-domain consistency, the algorithm rescues these attenuated paths using only if they spatially align with the detection in the reference band; otherwise, the candidate is discarded to preclude false alarms.
Fig. 3 shows an example of clustering (Rx19), where all CIR samples are plotted along the delay axis. The adaptation margins are set to dB and dB; the former was empirically optimized to balance noise suppression and signal retention, while the latter was determined based on the sensitivity analysis as the saturation point where the recovery of common paths stabilizes. The results demonstrate that multipath components are consistently detected across the two frequencies. It is noted that the resulting dynamic range (DR), defined as the difference between the peak power and the noise floor (), ranges from approximately to dB, while the spurious-free dynamic range (SFDR), defined as the difference between the peak power and the threshold level, ranges from to dB in the LoS scenario.
@@ IV-A Dominant Scattering Process
Figs. 6 depicts the spatial variation of the PDPs along the Rx trajectory for both scenarios. Specifically, Fig. 6(a) shows the relative PDPs, where the delay is normalized to the LoS delay, while Fig. 6(b) presents the normal PDPs as a function of the absolute delay. At both frequencies, the propagation characteristics exhibit pronounced sparsity and simplicity. In the LoS scenario, as shown in Fig. 6(a), two major scattering processes that vary consecutively along the Rx trajectory are identified as SB reflections from the north wall (NW) and south wall (SW), with their trajectories validated using the image method ray-tracing applied to the simplified model in Fig. 5(b). Aside from the three main components, such as the LoS path and the two SB reflections, only a few additional comonents are observed, mainly originating from double-bounce (DB) reflections (NWSW and SWNW) and other unidentified interactions with random objects. This phenomenon becomes more pronounced at GHz. In contrast, in the NLoS scenario, as shown in Fig. 6(b), two dominant propagation mechanisms that evolve consecutively along the Rx trajectory are approximately identified as the SB reflection from the SW and the DB reflection resulting from the SWNW interaction. Note that although the SWNW DB interaction can dominate in the NLoS example in Fig. 6(b), this work develops a Q-D model for the LoS scenario because the NLoS dataset is limited. The behavior of the extracted MPCs is illustrated in Figs. 7(a), 7(b), and 7(c), which present the variations in MPC distribution across the delay, AoD, and AoA domains, respectively, at both frequencies. In the LoS scenario, the SB reflections from the NW and SW dominate, whereas DB reflections are rarely observed, a trend that was further confirmed by the clustering results.
@@ IV-B Cluster Characteristics
Fig. 8(b) shows the path gain (PG) distribution of the clusters for each frequency band. The solid lines represent the LoS cluster PG across all Rx points in the LoS scenario, while the dashed lines denote the total PG of the NLoS clusters. For the LoS clusters, the average PG values are dB at GHz and dB at GHz, corresponding to a power difference of dB. This difference is dB larger than the additional FSPL of dB, indicating that the GHz band experiences slightly greater degradation. However, for the NLoS clusters, the average PG values are dB at GHz and dB at GHz, yielding a power difference of dB, which is dB larger than the additional FSPL.
@@
V Path Loss and LSP Characterization
Regardless of the specific channel model, channel parameters including PL and LSPs are essential for applying standard channel models such as 3GPP TR 38.901 [2]. In this section, the measured channels are characterized in terms of PL, DS, AS, and Rician -factor at both frequencies and in both scenarios.
V-A Path Loss